OpenStudy (anonymous):

Janet ran 3.8 km in 40 minutes.Work out her average speed in km/h. Answer is not important, rather the workings

OpenStudy (anonymous):

3.8*1/(2/3) = 3.8*3/2 = 1.9*3 =5.7 Janet ran at 5.7 km/h

OpenStudy (anonymous):

2/3=40/60?

OpenStudy (anonymous):

\[speed=distance/time\] 40 minutes=40/60=2/3 hours

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

@ mathematical pro

OpenStudy (anonymous):

the last question you helped me with was also finding the average speed, why is this question solved differently to the previous

OpenStudy (anonymous):

nope it was distance

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

Oh yes i mistaked ><

OpenStudy (anonymous):

http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e5e21ab0b8b1f45b4980983 this is also a finding average speed question which is solved different ly to this one

OpenStudy (anonymous):

is there a universal formula for finding average speed that always works regardless of the different units?

OpenStudy (anonymous):

distance / time to do this distance

OpenStudy (anonymous):

marleythemongoose : Average speed = (total distance covered( / (total time taken)???

OpenStudy (anonymous):

Here the correct answer his 3.8/(2/3) = 3.8*3/2 = 1.9*3 = 5.7 this is the same response that last time, but last time I calculed the distance Janet could have run in an hour with the same speed she ran to do the 3.8 km

OpenStudy (anonymous):

is it a different formula for different units of time??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

ok so for seconds you just divide distance by time

OpenStudy (anonymous):

and for minutes you do (minutes/ 1 hour) and the * the distance?

OpenStudy (anonymous):

............?

OpenStudy (anonymous):

Average speed = (total distance covered) / (total time taken) and you have the average spped in the unit : unit of the distance / unit of the time you just have to convert to havee the unit you wanbt

OpenStudy (anonymous):

sorry i just forgot to cilc post ....